# Spectral Structure of the Mixed Hessian of the Dispersionless Toda $\tau$-Function

### 24 Mar 2026 v1

@@ 5.7. Jacobi realization in the positive range
For general $p$, Proposition 5.16 yields only a signed representing measure for $\mathcal{G}_{p}$. From this point onward we restrict to the positive-measure range $1\leq p\leq s$. In that range, positivity is supplied by an external Hausdorff-moment theorem for Raney numbers. We record exactly the part of that input used here, in the normalization relevant for the present paper, before passing to the Jacobi/OPRL framework.
Lemma 5.18 (External Hausdorff-moment input in the present normalization).
@@
This lemma is the only place where the external Raney-moment positivity input is used. In what follows, we use only the existence of the positive measure and the support endpoint $\tau_{s}=\zeta_{c}^{-1}$; no explicit formula for the density is needed. These facts are supplied by [14, Thm. 5] together with [15]. For background on the corresponding Raney distributions, see also [7].
@@
In the notation of [14], this is exactly the Raney sequence $A_{n}(s,p)$. Therefore [14, Thm. 5], as corrected in [15], applies with the parameter pair $(p,r)=(s,p)$: the hypotheses are satisfied because $s\geq 1$ and $0\leq p\leq s$. It follows that $\{R_{s,p}(n)\}_{n\geq 0}$ is represented by a probability measure supported on $0,\tau_{s}]$, with
@@
The Jacobi realization suggests a further structural property of the compact remainder $\widetilde{C}^{(q)}_{*}$ introduced in Theorem 4.6. Let $\phi_{2}^{(q)},\phi_{3}^{(q)},\ldots$ denote the eigenvectors of $\widetilde{C}^{(q)}_{*,\perp}$, ordered by decreasing eigenvalue $\mu_{2,*}^{(q)}\geq\mu_{3,*}^{(q)}\geq\cdots$, and view each $\phi_{k}^{(q)}$ as a function on the lattice $\{p_{j}=q+js\}_{j\geq 0}$. Numerically, $\phi_{k}^{(q)}$ exhibits exactly k1k-1 sign changes in the index $j$ (Figure 4), a pattern stable across all values of $s$, $q$, and $\beta$ tested. This oscillation count is the hallmark of classical Sturm oscillation for self-adjoint Jacobi matrices [1, Ch. III]. For the range $1\leq p\leq s$, each scalar factor $\mathcal{G}_{p_{j}}$ admits such a Jacobi realization by Theorem 5.21. The observed nodal property of the full matrix $\widetilde{C}^{(q)}_{*}$ is therefore consistent with, though not a direct corollary of, the scalar Jacobi framework, since the Gram entries mix contributions from different Jacobi operators $J_{p_{j}}$. We leave a rigorous nodal theorem for $\widetilde{C}^{(q)}_{*}$ as an open problem.
The results of this section provide the analytic framework for the intermediate regime $\zeta_{c}<\zeta<\zeta_{\mathrm{univ}}$. The generating functions $\mathcal{G}_{p}(u)$ extend to the slit plane via their hypergeometric structure, have a resonant singularity of parametric excess $\gamma=2$ at the branch point, and admit a nonvanishing edge density; for $1\leq p\leq s$, $\mathcal{G}_{p}$ also has a Jacobi–Weyl realization. Thus the scalar spectral data persist beyond the loss of the weighted compact operator picture. Geometric singularity on the domain boundary occurs only at $\zeta=\zeta_{\mathrm{univ}}$: for $s\geq 3$ this is cusp formation, whereas for $s=2$ the critical map degenerates to the classical Joukowski segment. We treat this geometric threshold next in Section 6.
@@ 6. The geometric threshold $\zeta_{\mathrm{univ}}$
We first characterize the geometric transition itself. Below $\zeta_{\mathrm{univ}}$ the boundary is a smooth Jordan curve. At $\zeta_{\mathrm{univ}}$, the critical geometry depends on $s$: for $s\geq 3$ the boundary develops semicubical cusps, whereas for $s=2$ the Joukowski trace degenerates to a line segment. Above $\zeta_{\mathrm{univ}}$, the map is no longer injective; for $s\geq 3$ the boundary self-intersects, while for $s=2$ the unit-circle trace remains an ellipse although the exterior map ceases to be one-to-one. We then show that
@@
so the logarithmic spectral instability from Theorems 4.6 and 4.7 occurs while the conformal map is still univalent. Finally, we prove that the analytically continued scalar Gram data remain finite at $u=\zeta_{\mathrm{univ}}^{2}$. Thus the spectral singularity at $\zeta_{c}$ and the geometric singularity at $\zeta_{\mathrm{univ}}$ are genuinely distinct.
@@ 6.1. Geometric threshold
For the comparison with the analytic threshold $\zeta_{c}$, we only need the location of the geometric threshold at which the exterior map ceases to be univalent. The finer local description of the boundary singularity at the threshold is classical and will not be used below.
@@
At the critical value $\zeta=\zeta_{\mathrm{univ}}$, the boundary singularity is standard. For $s\geq 3$, the image of $|w|=1$ develops $s$ semicubical cusps; for $s=2$, the Joukowski trace degenerates to a line segment. These local descriptions are classical and are not needed for the spectral analysis below.
@@ 6.2. Separation of the analytic and geometric thresholds
6.3. Regularity of the continued scalar data at $\zeta_{\mathrm{univ}}$
We finally show that the continued scalar data do not develop any new singularity at the geometric threshold. This is the final step in separating analytic and geometric criticality.
Since $\zeta_{\mathrm{univ}}>\zeta_{c}$, the point
@@ 6.3. Regularity of the continued scalar data at $\zeta_{\mathrm{univ}}$
By Proposition 5.4, $\mathcal{G}_{p}$ satisfies the generalized hypergeometric differential equation in the variable $\xi=\zeta_{c}^{-2}u$. This equation is Fuchsian, and its only finite singular points are $\xi=0$ and $\xi=1$; equivalently, in the variable $u$ the only finite singular points are $u=0$ and $u=\zeta_{c}^{2}$; see [18, §16.8,§16.11]. Hence every $u_{0}>\zeta_{c}^{2}$ with $u_{0}\neq\zeta_{c}^{2}$ is an ordinary point of the differential equation.
@@
are finite for every $u>\zeta_{c}^{2}$. In particular, they are finite at
$u=\zeta_{\mathrm{univ}}^{2},$
and the discontinuity density satisfies
@@
The three regimes of the problem may now be summarized as follows: (i) For $0<\zeta<\zeta_{c}$, each symmetry sector is described by a compact weighted Gram operator with one stiff eigenvalue diverging logarithmically as $\zeta\uparrow\zeta_{c}$, while the remaining sectorial spectrum stays bounded. Equivalently, there are at most $s$ logarithmically diverging eigenvalues globally. (ii) For $\zeta_{c}<\zeta<\zeta_{\mathrm{univ}}$, the weighted operator realization breaks down, but the scalar data $\mathcal{G}_{p}$, $\sigma_{p}^{\mathrm{cont}}$, and $\rho_{p}$ admit a canonical analytic continuation, with Jacobi–Weyl interpretation for $1\leq p\leq s$. (iii) At $\zeta=\zeta_{\mathrm{univ}}$, the conformal map loses univalence; for $s\geq 3$ the boundary develops cusp singularities, whereas for $s=2$ the critical map degenerates to the Joukowski segment. In both cases, the continued scalar data remain finite.
@@ Conclusion
We have shown that for the $s$-fold symmetric one-harmonic polynomial family the first spectral instability of the mixed Hessian of the dispersionless Toda $\tau$-function is governed by analytic criticality of the inverse map, not by the later geometric loss of univalence of the conformal map. The square-root singularity of the inverse branch produces the borderline coefficient decay that drives the transition.
@@
On the subcritical side, for any fixed $\beta>0$, the weighted Hilbert realization of Definition 4.1 separates this mechanism into one stiff direction and a bounded soft sector in each symmetry block. Accordingly, in each such weighted realization each sector contains exactly one logarithmically diverging eigenvalue, while the remaining sectorial spectrum stays bounded; after removing the spike direction, each fixed soft eigenvalue converges to the corresponding eigenvalue of a compact limiting remainder. The intrinsic coefficient-level content is the existence of a single dominant singular direction in each sector from which these weighted spikes are realized. Beyond the analytic threshold, the weighted compact realization ceases to apply, but the scalar Gram data remain well defined after analytic continuation. They are described by generalized hypergeometric functions, admit a Cauchy–Stieltjes representation, and in the positive range fit naturally into the Jacobi/Weyl framework.
@@
The geometric threshold lies strictly beyond the analytic one. Consequently, the Toda Hessian becomes spectrally singular while the conformal map is still univalent and the boundary is still smooth, and the continued scalar data have finite upper and lower lateral boundary values at the univalence threshold. In this sense, analytic and geometric criticality are genuinely distinct in the present family.
@@
The results also admit a natural heuristic interpretation in adjacent frameworks. In the Laplacian-growth language, the rank-one logarithmic spike singles out a distinguished stiff deformation mode before any geometric singularity appears. In the matrix-model or potential-theoretic language, the proved statement is that the mixed second-order response concentrates onto a single dominant sectorial direction at criticality. These interpretive remarks are not used in the proofs and are included only to indicate the structural scope of the result.
@@
The argument isolates a structural mechanism that should persist beyond the one-harmonic leaf: a dominant branch point, borderline coefficient decay, and positive Gram structure. Several extensions therefore suggest themselves. For more general polynomial conformal maps one expects several competing dominant orbits and hence a finite-rank version of the present instability mechanism. It would also be natural to understand the bounded soft sector more conceptually, possibly through hidden Jacobi or total-positivity structure, and to sharpen the leading-order theory by deriving asymptotics for spectral gaps, sector dependence, and soft eigenvalues.
@@ Acknowledgments
The study was implemented in the framework of the Basic Research Program at HSE University in 2026.
@@ Appendix A Asymptotics of Raney numbers
This appendix provides explicit Stirling–type asymptotics for the Raney numbers in (2.6). These formulas are used in the Gram estimates and in Appendix D.
@@
Since $\zeta_{c}^{-1}=s^{s}/(s-1)^{s-1}$, this gives (A.1); squaring yields (A.2). ∎
@@
uniformly for $t\in[0,\theta]$. Here the terms of order $m$, the $\log m$ contribution, and the explicit $t$-dependent prefactor encoded in $A_{s,p}$ cancel identically; the remainder is uniform because all auxiliary functions of $t$ are smooth on the compact interval $0,\theta]$. Exponentiating yields
@@ Appendix B Rank–one tail extraction in the Gram blocks
This appendix records the coefficient–level structure behind the rank–one decomposition in Theorem 4.6. Throughout we fix a symmetry sector $q\in\{1,\dots,s\}$, write
@@ Appendix C Hypergeometric continuation of $\mathcal{G}_{p}$
We give a proof of the hypergeometric representation stated in Proposition 5.4. We keep the notation $\{\alpha_{i}\}_{i=1}^{2s}$ and $\{\beta_{j}\}_{j=1}^{2s-1}$ from (5.6).
@@ Appendix D The logarithmic coefficient at the branch point
This appendix computes the coefficient $B(\zeta_{c}^{2})$ in the resonant local expansion (5.7) of $\mathcal{G}_{p}$ at the branch point $u=\zeta_{c}^{2}$. Recall that the radius of convergence of $\mathcal{G}_{p}$ equals $\zeta_{c}^{2}$.

### 16 Apr 2026 v2

@@ 5.6. Jacobi realization in the positive range
For general $p$, Proposition 5.16 yields only a signed representing measure for $\mathcal{G}_{p}$. We now restrict to the positive range $1\leq p\leq s$, where positivity is supplied by an external Hausdorff-moment theorem for Raney numbers. Before passing to the Jacobi/OPRL framework, we state only the part needed here, in the normalization used in this paper.
Lemma 5.18 (External Hausdorff moment criterion in the present normalization).
@@
This is the only lemma that relies on positivity of the external Raney moments. In what follows, we use only the existence of the positive measure and the support endpoint $\tau_{s}=\zeta_{c}^{-1}$; no explicit formula for the density is needed. These facts are supplied by [13, Thm. 5] together with [14]. For background on the corresponding Raney distributions, see also [6].
@@
In the notation of [13], this is exactly the Raney sequence $A_{n}(s,p)$. Therefore [13, Thm. 5], as corrected in [14], applies with the parameter pair $(p,r)=(s,p)$: the hypotheses are satisfied because $s\geq 1$ and $0\leq p\leq s$. It follows that $\{R_{s,p}(n)\}_{n\geq 0}$ is represented by a probability measure supported on $0,\tau_{s}]$, with
@@
The Jacobi realization suggests a further structural property of the compact remainder $\widetilde{C}^{(q)}_{*}$ introduced in Theorem 4.6. Let $\phi_{2}^{(q)},\phi_{3}^{(q)},\ldots$ denote the eigenvectors of $\widetilde{C}^{(q)}_{*,\perp}$, ordered by decreasing eigenvalue $\mu_{2,*}^{(q)}\geq\mu_{3,*}^{(q)}\geq\cdots$, and view each $\phi_{k}^{(q)}$ as a function on the lattice $\{p_{j}=q+js\}_{j\geq 0}$. Numerically, $\phi_{k}^{(q)}$ exhibits exactly k1k-1 sign changes in the index $j$ (Figure 4), a pattern stable across all values of $s$, $q$, and $\beta$ tested. This oscillation count is the hallmark of classical Sturm oscillation for self-adjoint Jacobi matrices [1, Ch. III]. For the range $1\leq p\leq s$, each scalar factor $\mathcal{G}_{p_{j}}$ admits such a Jacobi realization by Theorem 5.21. The observed nodal property of the full matrix $\widetilde{C}^{(q)}_{*}$ is therefore consistent with the scalar Jacobi framework. It is not, however, a direct corollary of that framework, since the Gram entries combine contributions from different Jacobi operators $J_{p_{j}}$. We leave a rigorous nodal theorem for $\widetilde{C}^{(q)}_{*}$ as an open problem.
The results of this section provide the analytic framework for the intermediate regime $\zeta_{c}<\zeta<\zeta_{\mathrm{univ}}$. The generating functions $\mathcal{G}_{p}(u)$ extend to the slit plane via their hypergeometric structure, have a resonant singularity of parametric excess $\gamma=2$ at the branch point, and admit a nonvanishing edge density; for $1\leq p\leq s$, $\mathcal{G}_{p}$ also has a Jacobi–Weyl realization. Thus the scalar spectral quantities persist beyond the loss of the weighted compact operator picture. Geometric singularity on the domain boundary occurs only at $\zeta=\zeta_{\mathrm{univ}}$: for $s\geq 3$ this is cusp formation, whereas for $s=2$ the critical map degenerates to the classical Joukowski segment. We treat this geometric threshold next in Section 6.
@@ 6. The geometric threshold $\zeta_{\mathrm{univ}}$
We first characterize the geometric transition itself. Below $\zeta_{\mathrm{univ}}$ the boundary is a smooth Jordan curve. At $\zeta_{\mathrm{univ}}$, the critical geometry depends on $s$: for $s\geq 3$ the boundary develops semicubical cusps, whereas for $s=2$ the Joukowski trace degenerates to a line segment. Above $\zeta_{\mathrm{univ}}$, the map is no longer injective. For $s\geq 3$ the boundary self-intersects, while for $s=2$ the unit-circle trace remains an ellipse although the exterior map ceases to be one-to-one. We then show that
@@
so the logarithmic spectral instability from Theorems 4.6 and 4.7 occurs while the conformal map is still univalent. Finally, we prove that the analytically continued scalar Gram quantities remain finite at $u=\zeta_{\mathrm{univ}}^{2}$. Thus the spectral singularity at $\zeta_{c}$ and the geometric singularity at $\zeta_{\mathrm{univ}}$ are genuinely distinct.
@@ 6.1. Geometric threshold
For the comparison with the analytic threshold $\zeta_{c}$, we only need the location of the geometric threshold at which the exterior map ceases to be univalent. The finer local description of the boundary singularity at the threshold is classical.
@@
At the critical value $\zeta=\zeta_{\mathrm{univ}}$, the boundary singularity is standard. For $s\geq 3$, the image of $|w|=1$ develops $s$ semicubical cusps, while for $s=2$, the Joukowski trace degenerates to a line segment. These local descriptions are classical and are not needed for the spectral analysis below.
@@ 6.2. Separation of the analytic and geometric thresholds
6.3. Regularity of the continued scalar quantities at $\zeta_{\mathrm{univ}}$
We finally show that the continued scalar quantities do not develop any new singularity at the geometric threshold. This is the final step in separating analytic and geometric criticality. Since $\zeta_{\mathrm{univ}}>\zeta_{c}$, the point
@@ 6.3. Regularity of the continued scalar quantities at $\zeta_{\mathrm{univ}}$
By Proposition 5.4, $\mathcal{G}_{p}$ satisfies the generalized hypergeometric differential equation in the variable $\xi=\zeta_{c}^{-2}u$. This equation is Fuchsian, and its only finite singular points are $\xi=0$ and $\xi=1$. Equivalently, in the variable $u$ the only finite singular points are $u=0$ and $u=\zeta_{c}^{2}$; see [17, §16.8,§16.11]. Hence every $u_{0}>\zeta_{c}^{2}$ with $u_{0}\neq\zeta_{c}^{2}$ is an ordinary point of the differential equation.
@@
are finite for every $u>\zeta_{c}^{2}$. In particular, they are finite at $u=\zeta_{\mathrm{univ}}^{2}$, and the discontinuity density satisfies
@@
The three regimes of the problem may now be summarized as follows: (i) For $0<\zeta<\zeta_{c}$, each symmetry sector is described by a compact weighted Gram operator with one stiff eigenvalue diverging logarithmically as $\zeta\uparrow\zeta_{c}$, while the remaining sectorial spectrum stays bounded. Equivalently, there are at most $s$ logarithmically diverging eigenvalues globally. (ii) For $\zeta_{c}<\zeta<\zeta_{\mathrm{univ}}$, the weighted operator realization breaks down, but the scalar quantities $\mathcal{G}_{p}$, $\sigma_{p}^{\mathrm{cont}}$, and $\rho_{p}$ admit a canonical analytic continuation, with Jacobi–Weyl interpretation for $1\leq p\leq s$. (iii) At $\zeta=\zeta_{\mathrm{univ}}$, the conformal map loses univalence. For $s\geq 3$ the boundary develops cusp singularities, whereas for $s=2$ the critical map degenerates to the Joukowski segment. In both cases, the continued scalar quantities remain finite.
@@ Conclusion
We have shown that for the $s$-fold symmetric one-harmonic polynomial family the first spectral instability in the weighted block realizations of the mixed Hessian of the dispersionless Toda $\tau$-function is governed by analytic criticality of the inverse map, not by the later geometric loss of univalence of the conformal map. The square-root singularity of the inverse map produces the borderline coefficient decay that drives the transition.
@@
On the subcritical side, for any fixed $\beta>0$, the weighted Hilbert realization of Definition 4.1 separates this mechanism into one stiff direction and a bounded soft sector in each symmetry block. Accordingly, in each such weighted realization each sector contains exactly one logarithmically diverging eigenvalue, while the remaining sectorial spectrum stays bounded. After removing the spike direction, each fixed soft eigenvalue converges to the corresponding eigenvalue of a compact limiting remainder. At coefficient level, the same phenomenon is encoded by a single dominant singular direction in each sector, from which the weighted spikes are realized. Beyond the analytic threshold, the weighted compact realization ceases to apply, but the scalar Gram quantities remain well defined after analytic continuation. They are described by generalized hypergeometric functions, admit a Cauchy–Stieltjes representation, and in the positive range fit naturally into the Jacobi/Weyl framework.
@@
The geometric threshold lies strictly beyond the analytic one. Consequently, the weighted block theory becomes spectrally singular while the conformal map is still univalent and the boundary is still smooth, and the continued scalar quantities have finite upper and lower lateral boundary values at the univalence threshold. In this sense, analytic and geometric criticality are genuinely distinct in the present family.
@@
The result also has a direct interpretation in adjacent frameworks. In the Laplacian-growth language, the rank-one logarithmic spike singles out a distinguished stiff deformation mode before any geometric singularity appears. In the matrix-model or potential-theoretic language, the same statement says that the mixed second-order response concentrates onto a single dominant sectorial direction at criticality. These interpretations lie outside the proofs. They serve only to place the theorem in a broader geometric and physical context.
@@
The argument isolates a structural mechanism that should persist beyond the one-harmonic leaf: a dominant branch point, borderline coefficient decay, and positive Gram structure. For more general polynomial conformal maps, one expects several competing dominant orbits and hence a finite-rank version of the present instability mechanism. Another natural problem is to understand the bounded soft sector more conceptually, possibly through hidden Jacobi or total-positivity structure, and to sharpen the leading-order theory by deriving asymptotics for spectral gaps, sector dependence, and soft eigenvalues.
@@ Acknowledgments
This work is an output of the research project “Symmetry. Information. Chaos” implemented as a part of the Basic Research Program at National Research University Higher School of Economics (HSE University).
@@ Appendix A Asymptotics of Raney numbers
This appendix presents the Stirling-type asymptotics for the Raney numbers used in the Gram estimates of Section 4 and in the branch-point coefficient computation from Appendix D.
@@
Since $\zeta_{c}^{-1}=s^{s}/(s-1)^{s-1}$, this gives (A.1). Squaring yields (A.2). ∎
@@
uniformly for $t\in[0,\theta]$. Here the terms of order $m$, the $\log m$ contribution, and the explicit $t$-dependent prefactor encoded in $A_{s,p}$ cancel identically. The remainder is uniform because all auxiliary functions of $t$ are smooth on the compact interval $0,\theta]$. Exponentiating yields
@@ Appendix B Rank–one tail extraction in the Gram blocks
This appendix supplies the coefficient-level tail extraction used in the proof of Theorem 4.6. Throughout we fix a symmetry sector $q\in\{1,\dots,s\}$, write
@@ Appendix C Hypergeometric continuation of $\mathcal{G}_{p}$
This appendix proves the hypergeometric representation used in Section 5. We keep the notation $\{\alpha_{i}\}_{i=1}^{2s}$ and $\{\beta_{j}\}_{j=1}^{2s-1}$ from (5.6).
@@ Appendix D The logarithmic coefficient at the branch point
This appendix computes the coefficient $B(\zeta_{c}^{2})$ in the resonant local expansion (5.7) of $\mathcal{G}_{p}$ at the branch point $u=\zeta_{c}^{2}$. The argument isolates the universal $\operatorname{Li}_{3}$ tail in the coefficient asymptotics.
@@
Proof.
The first claim is immediate from the closed formula (D.1). The expression for $\rho_{p}(\zeta_{c}^{2})$ from Theorem 5.13 is a rational multiple of $B(\zeta_{c}^{2})$, so the same rational-over-$\pi$ structure is inherited by the edge value. ∎